# Beyond Next-Token Prediction: A Performance Characterization of Diffusion versus Autoregressive Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have achieved state-of-the-art performance on a broad range of Natural Language Processing (NLP) tasks, including document processing and coding. Autoregressive Language Models (ARMs), which generate tokens sequentially conditioned on all previous tokens, have been the predominant paradigm for LLMs. However, while these networks have achieved high accuracy across a range of downstream tasks, they exhibit low arithmetic intensity due to the inherent sequential dependency with next-token prediction. Recently, Diffusion Language Models (DLMs) have emerged as a promising alternative architecture. DLMs generate output text in parallel, breaking the limitations of sequential dependency. However, the performance implications of DLMs relative to commonly deployed ARMs are not fully understood. In this work, we present a comprehensive performance study analyzing the performance characteristics of ARMs and DLMs, using both theoretical analysis and profiling data to characterize the trade-offs between these approaches. We illustrate that although DLMs exhibit higher arithmetic intensity compared to ARMs because of their capability to utilize parallelism across sequence lengths, they fail to scale effectively to longer contexts. We then explore DLMs with block-wise decoding, outlining how this approach allows for increased arithmetic intensity, while still scaling well to long contexts (similar to ARMs). We also show interesting trade-offs for batched inference, where we find that ARMs exhibit superior throughput, as they benefit more from parallelism across sequences in the batch. Finally, we highlight opportunities for accelerating DLM inference, and, in particular, highlight the importance of reducing the number of sampling steps for allowing open-source DLMs to provide improved latency relative to ARMs.
@@ 1 Introduction
Large Language Models (LLMs) have ushered in a new era for machine learning and artificial intelligence, demonstrating broad applicability across instruction following, code generation, summarization, and multi-turn dialogue [24]. Modern LLMs are predominantly Autoregressive Language Models (ARMs): autoregressive, decoder-only Transformers trained with a next-token prediction objective under a causal attention mask [35, 8]. Although the training process is highly parallelizable thanks to teacher forcing, this objective imposes a strict sequential dependency at inference time. To mitigate this, various techniques have emerged, with speculative decoding being a prominent example that uses a smaller draft model to generate token bursts for parallel verification [18, 2, 20]. However, the drafting path remains sequential and offers limited speedups.
As a promising alternative, diffusion language models (DLMs) have recently gained attention for their new possibilities in text generation. Originating from the broader success of diffusion models in modalities like image, audio, and video synthesis [13, 28, 4, 9], DLMs operate by iteratively denoising a sequence of random noise into coherent text [21, 11, 29]. This process updates all token positions in parallel at each step, thereby breaking the token-level sequential dependency inherent to autoregressive decoding, and showing promise as an architecture that better exploits parallelism along the sequence dimension. Closed-source DLMs have publicly emphasized substantial speedups (e.g., Google’s experimental Gemini Diffusion [12] and Inception Labs’ Mercury [16] claim up to 10× faster and cheaper generation than strong autoregressive baselines), whereas current open-source DLMs often remain substantially slower than comparable ARMs in practice (see Section 3.3).
@@
  • Unified analysis. (Sections 3.2 and 3.3) We present profiling data along with roofline modeling analysis to characterize ARM prefill/decode and DLM variants across prompt and output lengths. Our analysis shows that DLMs have greater arithmetic intensity than ARMs during decoding; and also how DLMs scale poorly to long contexts, due to their need to re-process the full sequence during each sampling step.
  • Block-wise decoding and batching. (Sections 3.4 and 3.5) We extend the study to block-wise DLMs with approximate KV caching and to batched serving. We demonstrate how block-wise DLMs exhibit increased arithmetic intensity relative to ARMs, while still scaling well (similar to ARMs) to long context lengths. In contrast, for batched serving, ARMs exhibit superior throughput, as they better exploit sequence-level parallelism compared to DLMs (with or without block-wise diffusion).
@@ 2 Background
Autoregressive and Diffusion Language Models. Autoregressive language models (ARMs), the foundation of modern LLMs, are trained with a next-token prediction objective under a causal mask, restricting attention to past tokens [27, 1]. While training can be efficiently parallelized, inference is inherently sequential because each new token depends on the entire previously generated prefix [31]. This sequential dependency creates a performance bottleneck, leading to low arithmetic intensity and a decode path that is memory-bandwidth-bound [19, 17]. To avoid recomputing activations for past tokens at every step, KV caching has become standard. By storing Key (K) and Value (V) vectors during decoding, it largely eliminates recomputation and accelerates autoregressive inference [26]. The trade-off is a substantial increase in memory footprint and bandwidth pressure [14].
Diffusion models generate samples by initializing from Gaussian noise and iteratively applying reverse sampling steps, gradually transforming noise into a clean sample from the data distribution [13]. This family of models has been successfully adopted across modalities, including text-conditioned image generation [28], audio synthesis [4], video generation [9], and even language. In language modeling, diffusion language models (DLMs) generate over a fixed-length sequence via iterative refinement, updating all token positions in parallel at each step [21, 11, 29]. Architecturally, DLMs typically adopt Transformer architectures with bidirectional (non-causal) attention, allowing each position to attend to the full context [23]. Early work focused on continuous DLMs that operate in embedding space [10, 11, 29]; more recently, masked diffusion models (MDMs), a class of discrete DLMs that operate directly on categorical token spaces, have emerged as competitive alternatives to ARMs [34, 23]. A key systems challenge for DLMs is that bidirectional attention is incompatible with standard KV caching, causing latency to increase substantially as context length grows (see Section 3.3). To mitigate this, block-wise decoding divides the sequence into blocks. Decoding proceeds in an autoregressive manner across blocks, while positions within each block are updated in parallel, enabling approximate block-level KV caching [33, 22, 15]. A separate line of work, speculative diffusion decoding, keeps the main model autoregressive while employing a diffusion model as the drafter [6].
@@
One method to visualize bottlenecks across prompt lengths, KV caching, and batch sizes is the roofline model [32]. The roofline model provides a visual criterion for classifying programs as compute-bound or memory-bound via the hardware ceilings for compute and bandwidth. The ridge point is the knee where the bandwidth roof meets the compute roof, occurring at $\mathrm{AI}_{\text{ridge}}=\frac{\text{peak compute}}{\text{peak memory bandwidth}}$ (FLOP/byte); workloads with $\mathrm{AI}<\mathrm{AI}_{\text{ridge}}$ are memory-bound, and those with $\mathrm{AI}\geq\mathrm{AI}_{\text{ridge}}$ are compute-bound.
@@ 3.1 Experimental Setup
Models. We study the LLaMA-3-8B-Instruct model [8] (an ARM) and the LLaDA-8B-Instruct model [23] (a DLM) as representative baselines (as they have comparable model size and comparable downstream task performance as reported in [23]). For block-wise DLM experiments with approximate KV caching (outlined in Section 3.4), we adopt the LLaDA variant (dual cache) from Fast-dLLM [33], as the original LLaDA does not support approximate KV caching.
Methodology. We extend prior analytical modeling frameworks [17, 30] to (i) decompose ARM inference into prefill and decode, and (ii) model DLM inference as multi-step refinement with optional block-wise KV reuse.
@@ 3.2 Asymptotic Analysis of Arithmetic Intensity in ARM and DLM
ARMs. Prefill shifts from linear-projection-dominated AI ($\mathcal{O}(BL_{p})$) when $L_{p}\ll d$ to attention-dominated AI ($\mathcal{O}(L_{p})$) when $L_{p}\gg d$. During the subsequent decode phase, each step involves small matrix multiplications (linear projections) and retrieval of the growing KV cache for attention. As a result, the AI is initially $\mathcal{O}(B)$ when $L_{p}\ll d$, but it drops to $\mathcal{O}(1)$ once $L_{p}\gg d$, since the memory load from attention begins to dominate the fixed per-token compute; see also [30].
Block-DLMs. In a block-wise DLM with block size $G$, each refinement step updates a block of $G$ tokens in parallel. The parallel computation within its linear layers is analogous to an ARM prefill operation on a sequence of length $G$. However, the attention mechanism is fundamentally distinct, as these G tokens attend to the KV cache representing the entire sequence $L$. The one-time cost of initializing this cache via a full-attention pass is omitted, as it is a less dominant term compared to the cumulative cost of the $K$ refinement steps.
@@ 3.4 Implications of Block-wise Decoding with KV Cache
Block-wise decoding combines the strengths of ARMs and DLMs: it reuses the KV cache across completed blocks (ARM-like), while updating all tokens within the active block in parallel (DLM-like). The diffusion steps $K$ are typically distributed evenly across blocks. We conduct a profiled analysis, where we fix the prompt length to $L_{p}=1\mathrm{k}$, consider block sizes $G\in\{32,64,128\}$, and set the number of diffusion steps equal to the generation length ($K=L_{g}$). Figure 3 (a) compares end-to-end latency versus generation length $L_{g}$, demonstrating how block-wise decoding with KV caching reduces latency by roughly $2\times$ to $3\times$, compared to naive DLM (labeled "No Blocks"). Moreover, Figure 3 (b) shows that this approach not only reduces the AI but also makes it invariant to the generation length, depending only on the block size $G$. This reduction in AI translates into lower latency in compute-bound settings (e.g., naive DLM inference in Fig. 3 (b)), and block-wise decoding reduces the required FLOPs relative to naive DLM (as highlited in Table 1).
Note that even with KV caching, the DLM remains slower than the ARM at matched prompt and output lengths when $K$ scales with $L_{g}$. For instance, the widely used configuration of $K=L_{g}$ [23, 34] implies that the model denoises and finalizes only one token per step on average, which is hugely inefficient. This underscores the importance of reducing diffusion steps via efficient multi-token denoising, a point we revisit in Section 4.
@@ 3.5 Implications of Inference with Different Batch Sizes
Real-world applications often serve multiple users concurrently, making scalability with respect to the batch size $B$ a critical metric. We therefore measure throughput (tokens/s) as a function of $B$, fixing the generation length at $L_{g}=128$ and evaluating two prompt categories: short ($L_{p}=128$) and long ($L_{p}=2\text{k}$). We set the block size to $G=32$ and use $K=L_{g}$ for the block-wise DLM. All experiments in this section are run on a single NVIDIA A100 (80 GB) GPU to accommodate the memory demands of batched inference. We omit the naive DLM, as prior work shows it does not benefit from batching due to its consistently compute-bound bidirectional full attention [33].
@@ 4 Discussion and Future Work
Our analysis shows that a naive DLM exhibits substantial latency overheads relative to ARMs as the context length increases. To ameliorate this, block-wise decoding with approximate KV caching offers an effective hybrid approach. This method achieves increased arithmetic intensity (AI) by exploiting intra-block parallelism, while still scaling well to long contexts, similar to ARMs. However, in batched serving, the block-wise DLM’s parallel generation saturates the GPU at smaller batch sizes than ARMs. Combined with the DLM’s high cost of iterative refinement, this results in ARMs providing distinct throughput advantages in the large-batch regime.
We also observe that open-source DLM implementations typically lag behind open-source ARMs in speed. In contrast, recent closed-source systems [12, 25, 16] report striking latency gains at comparable quality – often hundreds to thousands of times faster than current open-source DLMs, and even faster than ARMs. Two core bottlenecks hinder open-source DLM performance: poor scaling with context length and high refinement costs.
For the scaling problem, Section 3.4 confirms that block-wise decoding with approximate KV caching [33, 22, 15] is an indispensable strategy. Yet, to truly rival ARM speed, the most critical path forward involves reducing the number of sampling steps $K$. In common configurations, $K$ scales with the generation length $L_{g}$, leading to high per-output costs and compute saturation, as highlighted in Section 3.5. Recent approaches attempt to cut steps without quality loss, such as multi-token finalization with autoregressive guidance [15], confidence-based early finalization [33], and distillation into few-step or even one-step student models [5]. Reducing sampling steps is particularly promising in small-batch scenarios, where a DLM’s throughput could surpass an ARM’s, as mentioned in Appendix B. Finally, orthogonal system-level optimizations can offer additional speedups. These include compute-focused methods like low-precision execution as well as sparse attention methods that utilize only the most salient cache entries.
@@ Appendix A RTX A6000 Roofline Parameters
Appendix B Implications of Reduction in Sampling Steps

### v2

@@ Abstract
Large Language Models (LLMs) have achieved state-of-the-art performance on a broad range of Natural Language Processing (NLP) tasks, including document processing and code generation. Autoregressive Language Models (ARMs), which generate tokens sequentially conditioned on all previous tokens, have been the predominant paradigm for LLMs. While these models have achieved high accuracy across a range of downstream tasks, they exhibit low arithmetic intensity due to the inherent sequential dependency in next-token prediction. Recently, Diffusion Language Models (DLMs) have emerged as a promising alternative architecture. DLMs generate output tokens in parallel, mitigating the limitations of sequential decoding. However, the performance implications of DLMs relative to commonly deployed ARMs are not fully understood. In this work, we present a comprehensive study of the performance characteristics of ARMs and DLMs, combining theoretical analysis with empirical profiling to characterize the trade-offs between these approaches. We show that although DLMs can achieve higher arithmetic intensity than ARMs by leveraging parallelism across token positions, they fail to scale effectively with longer contexts. We then explore block-wise decoding for DLMs, which decouples arithmetic intensity from sequence length and enables better scaling to long contexts (similar to ARMs). We also examine batched inference and find that ARMs exhibit superior throughput as they benefit more from parallelism across sequences in the batch. Finally, we highlight opportunities for accelerating DLM inference, emphasizing that reducing the number of sampling steps is key for open-source DLMs to achieve lower latency relative to ARMs.
@@ 1 Introduction
Large Language Models (LLMs) have ushered in a new era for machine learning and artificial intelligence, demonstrating broad applicability across instruction following, code generation, summarization, and multi-turn dialogue [openai2023gpt4]. Modern LLMs are predominantly Autoregressive Language Models (ARMs): decoder-only Transformers trained with a next-token prediction objective under a causal attention mask [zhang2022opt, dubey2024llama]. Although the training process is highly parallelizable thanks to teacher forcing, this objective imposes a strict sequential dependency at inference time. To mitigate this, various techniques have emerged, with speculative decoding being a prominent example that uses a smaller draft model to generate token bursts for parallel verification [kim2023speculative, chen2023accelerating, leviathan2023fast]. However, the drafting path remains sequential and offers limited speedups.
[…]
@@
  • Unified analysis. (Sections 3.2 and 3.3) We present profiling data along with roofline modeling analysis to characterize ARM prefill/decode and DLM variants across prompt and output lengths. Our analysis shows that DLMs have greater arithmetic intensity (AI) than ARMs during decoding, but scale poorly to long contexts due to their need to re-process the full sequence at each sampling step.
  • Block-wise decoding and batching. (Sections 3.4 and 3.5) We demonstrate how block-wise DLMs reduce AI relative to naive DLM decoding, while maintaining higher AI than ARMs and scaling well to long context lengths. In contrast, for batched serving, ARMs exhibit superior throughput, as they better exploit sequence-level parallelism compared to DLMs (with or without block-wise decoding).
@@ 2 Background
Autoregressive and Diffusion Language Models. Autoregressive language models (ARMs), the foundation of modern LLMs, are trained with a next-token prediction objective under a causal mask, restricting attention to past tokens [radford2019language, brown2020language]. While training can be efficiently parallelized, inference is inherently sequential because each new token depends on the entire previously generated prefix [vaswani2017attention]. This sequential dependency creates a performance bottleneck, leading to low arithmetic intensity and a decode path that is memory-bandwidth-bound [kim2023squeezellm, kim2023full]. To avoid recomputing activations for past tokens at every step, KV caching has become standard. By storing Key (K) and Value (V) vectors during decoding, it largely eliminates recomputation and accelerates autoregressive inference [pope2022efficient]. The trade-off is a substantial increase in memory footprint and bandwidth pressure [hooper2024kvquant].
Diffusion models generate samples by initializing from Gaussian noise and iteratively applying reverse sampling steps, gradually transforming noise into a clean sample from the data distribution [ho2020ddpm]. This family of models has been successfully adopted across modalities, including text-conditioned image generation [ramesh2022hierarchical], audio synthesis [chen2021wavegrad], video generation [esser2023sceneguided], and even language. In language modeling, diffusion language models (DLMs) generate over a fixed-length sequence via iterative refinement, updating all token positions in parallel at each step [li2022diffusionlm, gong2023diffuseq, strudel2022sed]. Architecturally, DLMs typically adopt Transformer architectures with bidirectional (non-causal) attention, allowing each position to attend to the full context [nie2025large]. Early work focused on continuous DLMs that operate in embedding space [gao2024embedding, gong2023diffuseq, strudel2022sed]; more recently, masked diffusion models (MDMs), a class of discrete DLMs that operate directly on categorical token spaces, have emerged as competitive alternatives to ARMs [dream2025, nie2025large]. A key systems challenge for DLMs is that bidirectional attention is incompatible with standard KV caching, causing latency to increase substantially as context length grows (see Section 3.3). To mitigate this, block-wise decoding divides the sequence into blocks. Decoding proceeds in an autoregressive manner across blocks, while positions within each block are updated in parallel, enabling approximate KV caching of the prompt and non-active blocks. [wu2025fastdllm, ma2025dkvcache, hu2025freecache]. A separate line of work, speculative diffusion decoding, keeps the main model autoregressive while employing a diffusion model as the drafter [christopher2024specdiff].
@@
One method to visualize bottlenecks across prompt lengths, KV caching, and batch sizes is the roofline model [williams2009roofline]. The roofline model provides a visual criterion for classifying programs as compute-bound or memory-bound via the hardware ceilings for compute and bandwidth. The ridge point is the knee where the bandwidth roof meets the compute roof, occurring at $\mathrm{AI}_{\text{ridge}}=\frac{\text{peak compute}}{\text{peak memory bandwidth}}$ (FLOP/byte); workloads with $\mathrm{AI}<\mathrm{AI}_{\text{ridge}}$ are memory-bound, and those with $\mathrm{AI}\geq\mathrm{AI}_{\text{ridge}}$ are compute-bound.
@@ 3.1 Experimental Setup
Models. We study the LLaMA-3-8B-Instruct model [dubey2024llama] (an ARM) and the LLaDA-8B-Instruct model [nie2025large] (a DLM) as representative baselines (as they have comparable model size and comparable downstream task performance as reported in [nie2025large]). For block-wise DLM experiments with approximate KV caching (outlined in Section 3.4), we adopt the LLaDA variant (dual cache) from Fast-dLLM [wu2025fastdllm], as the original LLaDA does not support KV caching.
Methodology. We extend prior analytical modeling frameworks [kim2023full, tiwari2025quantspec] to (i) decompose ARM inference into prefill and decode, and (ii) model DLM inference as multi-step refinement with optional block-wise KV reuse.
@@ 3.2 Asymptotic Analysis of Arithmetic Intensity in ARM and DLM
ARMs. Prefill shifts from linear-projection-dominated AI ($\mathcal{O}(BL_{p})$) when $L_{p}\ll d$ to attention-dominated AI ($\mathcal{O}(L_{p})$) when $L_{p}\gg d$. During the subsequent decode phase, each step involves small matrix multiplications (linear projections) and retrieval of the growing KV cache for attention. As a result, the AI is initially $\mathcal{O}(B)$ when $L_{p}\ll d$, but it drops to $\mathcal{O}(1)$ once $L_{p}\gg d$, since the memory load from attention begins to dominate the fixed per-token compute; see also [tiwari2025quantspec].
Block-DLMs. In a block-wise DLM with block size $G$, each refinement step updates a block of $G$ tokens in parallel. The parallel computation within its linear layers is analogous to an ARM prefill operation on a sequence of length $G$. However, the attention mechanism is fundamentally distinct, as these G tokens attend to the KV cache representing the entire sequence $L$. The cost of initializing and periodically updating this cache via a full-attention pass is omitted, as it is a less dominant term compared to the cumulative cost of the $K$ refinement steps.
@@ 3.4 Implications of Block-wise Decoding with KV Cache
Block-wise decoding combines the strengths of ARMs and DLMs: it reuses the KV cache for the prompt and non-active blocks (ARM-like), while updating all tokens within the active block in parallel (DLM-like). The diffusion steps $K$ are typically distributed evenly across blocks. We conduct a profiled analysis, where we fix the prompt length to $L_{p}=1\mathrm{k}$, consider block sizes $G\in\{32,64,128\}$, and set the number of diffusion steps equal to the generation length ($K=L_{g}$). Figure 3 (a) compares end-to-end latency versus generation length $L_{g}$, demonstrating how block-wise decoding with KV caching reduces latency by roughly $2\times$ to $3\times$, compared to naive DLM (labeled "No Blocks"). Moreover, Figure 3 (b) shows that this approach not only reduces the AI but also makes it invariant to the generation length, depending only on the block size $G$. This reduction in AI translates into lower latency in compute-bound settings (e.g., naive DLM inference in Fig. 3 (b)), and block-wise decoding reduces the required FLOPs relative to naive DLM (as highlighted in Table 1). Furthermore, block-wise decoding preserves (and can even improve) task accuracy despite these efficiency gains; we report accuracy results in Appendix B.
Note that even with KV caching, the DLM remains slower than the ARM at matched prompt and output lengths when $K$ scales with $L_{g}$. For instance, the widely used configuration of $K=L_{g}$ [nie2025large, dream2025] implies that the model denoises and finalizes only one token per step on average, which is hugely inefficient. This underscores the importance of reducing diffusion steps via efficient multi-token denoising, a point we revisit in Section 4.
@@ 3.5 Implications of Inference with Different Batch Sizes
Real-world applications often serve multiple users concurrently, making scalability with respect to the batch size $B$ a critical metric. We therefore measure throughput (tokens/s) as a function of $B$, fixing the generation length at $L_{g}=128$ and evaluating two prompt categories: short ($L_{p}=128$) and long ($L_{p}=2\mathrm{k}$). We set the block size to $G=32$ and use $K=L_{g}$ for the block-wise DLM. All experiments in this section are run on a single NVIDIA A100 (80 GB) GPU to accommodate the memory demands of batched inference. We omit the naive DLM, as prior work shows it does not benefit from batching due to its consistently compute-bound bidirectional full attention [wu2025fastdllm].
@@ 4 Discussion and Future Work
Our analysis shows that a naive DLM exhibits substantial latency overheads relative to ARMs as the context length increases. To ameliorate this, block-wise decoding with approximate KV caching offers an effective hybrid approach. This method achieves increased arithmetic intensity (AI) by exploiting intra-block parallelism, while still scaling well to long contexts, similar to ARMs. However, in batched serving, the block-wise DLM’s parallel generation saturates the GPU at smaller batch sizes than ARMs. Combined with the DLM’s high cost of iterative refinement, this results in ARMs providing distinct throughput advantages in the large-batch regime. Recent closed-source systems [deepmind2025geminidiffusion, inceptionlabs2025bedrock, khanna2025mercury] report striking latency gains at comparable quality, even faster than ARMs. Based on our analysis, we therefore conclude that two core bottlenecks are responsible for the slowdown of current open-source DLMs: poor scaling with context length and high refinement costs.
For the scaling problem, Section 3.4 confirms that block-wise decoding with approximate KV caching [wu2025fastdllm, ma2025dkvcache, hu2025freecache] is an indispensable strategy. Yet, to truly rival ARM speed, the most critical path forward involves reducing the number of sampling steps $K$. In common configurations, $K$ scales with the generation length $L_{g}$, leading to high per-output costs and compute saturation, as highlighted in Section 3.5. Recent approaches attempt to cut steps without quality loss, such as multi-token finalization with autoregressive guidance [hu2025freecache], confidence-based early finalization [wu2025fastdllm], and distillation into few-step or even one-step student models [chen2025dlmone]. Reducing sampling steps is particularly promising in small-batch scenarios, where a DLM’s throughput could surpass an ARM’s, as mentioned in Appendix C. Finally, orthogonal system-level optimizations can offer additional speedups. These include compute-focused methods like low-precision execution as well as sparse attention methods that utilize only the most salient cache entries.
@@ Appendix A RTX A6000 Roofline Parameters
Appendix B Impact of Block-wise Decoding on Generation Quality
Table 2: DLM performance on GSM8K and HumanEval with and without block-wise decoding. We report results for block sizes $G\in\{64,32,16,8\}$ and the non-block-wise baseline (None). The generation length is fixed to $L_{g}=128$. GSM8K is reported as exact-match (EM) accuracy, and HumanEval is reported as pass@1.
[…]
None 63.2 13.4
64 73.8 34.1
32 73.6 35.4
16 75.7 38.4
8 75.0 38.4
As discussed in Section 3.4, block-wise decoding combined with KV caching provides systematic efficiency benefits by reducing redundant computation, thereby lowering latency under compute-bound DLM inference. Following the evaluation protocol of prior work [wu2025fastdllm], we evaluate block-wise DLMs on representative benchmarks, including the GSM8K math benchmark [cobbe2021gsm8k] and the HumanEval coding benchmark [chen2021evaluating].
As shown in Table 2, block-wise decoding consistently improves accuracy across both math and coding tasks compared to the naive DLM baseline (None). While the block size $G$ is a hyperparameter whose optimal value may vary across tasks, all block-wise configurations outperform the naive DLM in our experiments.
Two factors likely contribute to this behavior. First, block-wise decoding mitigates the excessive generation of end-of-text (|EOT|) tokens. In naive DLMs, generation can collapse because the model can "cheat" by repeatedly filling |EOT| from later positions, since this is easy, leading to degenerate outputs. Block-wise decoding helps alleviate this issue by enforcing parallel generation within the active block. Second, block-wise decoding introduces a weak left-to-right inductive bias at the block level. This ordering constraint can be beneficial for reasoning tasks, where a left-to-right structure naturally aligns with step-by-step reasoning processes.
Appendix C Implications of Reduction in Sampling Steps